In a certain town the annual real - estate tax on a property assessed at $121800 is $748. Find the tax due on a property assessed at $90000. Round your answer to two decimal places.
Select the correct answer.
a. $552.71
b. $526.47
c. $565.83
d. $583.84

Answer by TimothyLamb(4379)

748/121800 = t/90000
(748/121800)90000 = t
t = $552.71
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php